OpenStudy (anonymous):

3/x+2 +2/x-2=8/(x+2)(x-2) Solve and tell if its extraneous

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

now you can have extraneous solutions because you have denominators

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[\frac{x}{x+2}+\frac{2}{x-2}=\frac{8}{(x+2)(x-2)}\] you have a choice, you can either add on the left, or multiply both sides by \((x+2)(x-2)\) to clear the fractions

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

ok then the \((x+2)(x-2)\) cancels on the right leaving \(8\)

OpenStudy (anonymous):

on the left the first term will be \(x(x-2)\) and the second term will be \(2(x+2)\)

OpenStudy (anonymous):

by cancelling that leaves you with the equation \[x(x-2)+2(x+2)=8\] multiply out and solve the resulting quadratic

OpenStudy (anonymous):

x^2-2x+2x+4=8 x^2+4=8 x^2=4

OpenStudy (anonymous):

ok and if \(x^2=4\) then \(x=?\)

OpenStudy (anonymous):

x=+ or - 2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and in this case

OpenStudy (anonymous):

BOTH are extraneous, because both values would make the denominator equal to zero

OpenStudy (anonymous):

:) Thanks!

OpenStudy (anonymous):

yw